# TabNSA: Native Sparse Attention for Efficient Tabular Data Learning thanks: Code on GitHub: https://github.com/aseslamian/TabNSA

### v1

@@ IV-B Preprocessing
In order to handle missing values and class imbalance we used the preprocessing method in [13]. Let the input variable space as $\mathcal{D}ata\in\{\mathbb{R}\cup\mathbb{C}\cup\mathbb{B}\cup\varnothing\}$, where $\mathbb{R}$, $\mathbb{C}$, $\mathbb{B}$ and $\varnothing$ represent the domains of numerical, categorical, binary dat and NULL values respectively. After preprocessing block, we have $\{\mathcal{X},\mathcal{Y}\}$ denote a feature-target pair, where $x$ represents numerical features, and $y$ is a continuous variable for regression tasks or a whole number for classification tasks.
IV-C Experiment Set up
@@
To enhance TabNSA’s performance and stability, we systematically tuned critical hyperparameters using OPTUNA [20]. This involved dividing each dataset into training (70%), validation (10%), and test (20%) sets, training TabNSA on the training set, and evaluating it on the validation set to guide the hyperparameter optimization process.
@@
We employed the OPTUNA package to optimize key parameters, ensuring an efficient and automated search for the best configuration. Hyperparameter optimization is crucial for tabular data, given the variability in feature distributions and dataset characteristics. Once the optimal parameters are determined, the final model is tuned on train dataset and evaluated on the test set using standard metrics to assess its generalization performance. The search space included architectural parameters such as head dimension (8-46), number of attention heads (1-8), sliding window size (1-8), compression block size (4-16), selection block size (2-compression block size), and number of selected blocks (1-4). Furthermore, we fine-tuned optimization parameters like learning rate (1e-4 - 1e-3) and batch size (32-128) to strike a balance between convergence speed and generalization.
IV-D Binary Classification
@@
We evaluated TabNSA across eight datasets for the supervised binary classification task, with the results summarized in Table II. We include the results of other methods directly from [15]. To ensure a fair comparison, we followed the instructions in [15], ran the model 10 times, and reported the average. The model consistently outperformed all competing approaches, including classical machine learning methods, deep neural networks (DNNs), transformer-based architectures, and tree-based models.
Notably, TabNSA ranked first across all datasets, demonstrating its effectiveness. While it may not achieve the highest performance on every dataset, it consistently excels overall.
TABLE II: Performance Comparison of Models for Binary Classification
@@
Logistic Regression 0.720 0.836 0.557 0.851 0.748 0.801 0.769 0.860 11.5 (2.33)
XGBoost 0.726 0.895 0.587 0.912 0.892 0.821 0.758 0.925 6.69 (4.05)
MLP 0.643 0.832 0.568 0.904 0.613 0.832 0.779 0.893 10.9 (2.01)
SNN 0.641 0.880 0.540 0.902 0.621 0.834 0.794 0.892 9.81 (3.66)
TabNet 0.585 0.800 0.478 0.904 0.680 0.819 0.742 0.896 12.7 (1.67)
DCN 0.739 0.870 0.674 0.913 0.848 0.840 0.768 0.915 5.5 (2.41)
AutoInt 0.744 0.866 0.672 0.913 0.808 0.844 0.762 0.916 5.5 (2.83)
TabTrans 0.718 0.860 0.648 0.914 0.855 0.820 0.794 0.882 7.56 (4.12)
FT-Trans 0.739 0.859 0.657 0.913 0.862 0.841 0.793 0.915 5.5 (2.46)
VIME 0.735 0.852 0.485 0.912 0.769 0.837 0.786 0.908 8.31 (2.55)
SCARF 0.733 0.861 0.663 0.911 0.719 0.833 0.758 0.919 7.88 (2.89)
TransTab 0.768 0.881 0.643 0.907 0.851 0.845 0.822 0.919 4.56 (2.61)
TabMixer 0.660 0.907 0.659 0.900 0.829 0.821 0.974 0.969 6.19 (4.80)
TabNSA 0.874 0.898 0.717 0.967 0.721 0.926 0.979 0.967 2.38 (3.11)
IV-E Multi-class Classification
Table III presents a performance comparison of TabNSA and other neural network models on two multi-class classification datasets. Since multi-class classification tasks often suffer from class imbalance, we used macro-F1 as the primary evaluation metric during training to ensure a balanced assessment across all classes, following [21]. The results show that TabNSA consistently outperforms all models on both datasets.
TABLE III: Comparison of different methods on SG and FO datasets.
@@ IV-E Multi-class Classification
MLP TabTrans TabNet TabNSA

### v2

@@ 5.2 Preprocessing
In order to handle missing values and class imbalance we used the preprocessing method in [7]. We used this setting for Supervised and transfer learning task. Let the input variable space as $\mathcal{D}ata\in\{\mathbb{R}\cup\mathbb{C}\cup\mathbb{B}\cup\varnothing\}$, where $\mathbb{R}$, $\mathbb{C}$, $\mathbb{B}$ and $\varnothing$ represent the domains of numerical, categorical, binary dat and NULL values respectively. After preprocessing block, we have $\{\mathcal{X},\mathcal{Y}\}$ denote a feature-target pair, where $x$ represents numerical features, and $y$ is a continuous variable for regression tasks or a whole number for classification tasks.
5.3 Experiment Set up
@@
To enhance TabNSA’s performance and stability, we employed OPTUNA [29] to systematically tune critical hyperparameters through an efficient and automated search. Given the variability in feature distributions and dataset characteristics typical of tabular data, hyperparameter optimization is essential. Each dataset was divided into training (70%), validation (10%), and test (20%) sets, with the training and validation subsets used to guide the optimization process.
@@
In our experimental setup, we employed a two-stage optimization strategy. In the first stage, we utilized the OPTUNA package to automate the hyperparameter search specifically targeting the architecture of the NSA model. The search space included architectural parameters such as head dimension (8-46), number of attention heads (1-8), sliding window size (1-8), compression block size (4-16), selection block size (2-compression block size), and number of selected blocks (1-4). Furthermore, we fine-tuned optimization parameters like learning rate (1e-4 - 1e-3) and batch size (32-128) to strike a balance between convergence speed and generalization.
In the second stage, having fixed the optimal architectural configuration, we trained the final model’s MLP layers using the AdamW optimizer with a CrossEntropy loss function, finely tuning the learning rate ($1\times 10^{-4}$ to $1\times 10^{-3}$) and batch size (32–128) to balance convergence speed and generalization. The model was first transferred to the designated device (CPU or GPU) and then trained in mini-batches: each batch underwent a forward pass to compute predictions, followed by loss calculation, backpropagation, and parameter updates. If a validation loader was provided, the model was evaluated at the end of each epoch without gradient tracking, and the average validation loss was monitored to implement early stopping—halting training when no improvement in validation loss was observed for a predefined number of epochs and restoring the best-performing weights. Finally, the model’s performance was assessed on the test set using standard metrics, and a history of training and validation losses was returned to facilitate analysis of convergence and generalization.
For each trial, Optuna suggests values for both the architectural and optimization hyperparameters of the NSA model. This first-stage optimization integrates the full training procedure, including second-stage weight updates, to identify the hyperparameter configuration that maximizes validation AUC. Once the optimal parameters are identified, the final model is retrained on the training set and evaluated on the test set using standard metrics to assess generalization performance.
5.4 Binary Classification
@@
We evaluated TabNSA on eight datasets for supervised binary classification, with results summarized in Table 3. Baseline results from other methods are taken directly from [16]. Following their evaluation protocol, we ran TabNSA 10 times with different seeds and reported the average performance. TabNSA consistently outperformed classical machine learning methods, deep neural networks (DNNs), Transformer-based architectures, and tree-based models. Although it may not achieve the best score on every dataset, TabNSA ranked first overall, demonstrating strong and consistent performance.
To assess the contribution of its components, we compare TabNSA with the NSA module alone and with TabMixer in its default configuration from [7]. While NSA alone achieves notable improvements, TabNSA achieves a lower average rank (1.62 vs. 2.88) and a smaller standard deviation (0.744 vs. 3.36), indicating superior overall performance and greater stability across datasets.
Table 3: Performance Comparison of Models for Binary Classification
@@
Logistic Regression 0.720 0.836 0.557 0.851 0.748 0.801 0.769 0.860 12.5 (2.33)
XGBoost 0.726 0.895 0.587 0.912 0.892 0.821 0.758 0.925 7.69 (4.05)
MLP 0.643 0.832 0.568 0.904 0.613 0.832 0.779 0.893 11.9 (2.01)
SNN 0.641 0.880 0.540 0.902 0.621 0.834 0.794 0.892 10.8 (3.66)
TabNet 0.585 0.800 0.478 0.904 0.680 0.819 0.742 0.896 13.7 (1.67)
DCN 0.739 0.870 0.674 0.913 0.848 0.840 0.768 0.915 6.5 (2.41)
AutoInt 0.744 0.866 0.672 0.913 0.808 0.844 0.762 0.916 6.5 (2.83)
TabTrans 0.718 0.860 0.648 0.914 0.855 0.820 0.794 0.882 8.56 (4.12)
FT-Trans 0.739 0.859 0.657 0.913 0.862 0.841 0.793 0.915 6.5 (2.46)
VIME 0.735 0.852 0.485 0.912 0.769 0.837 0.786 0.908 9.31 (2.55)
SCARF 0.733 0.861 0.663 0.911 0.719 0.833 0.758 0.919 8.88 (2.89)
TransTab 0.768 0.881 0.643 0.907 0.851 0.845 0.822 0.919 5.56 (2.61)
TabMixer † 0.660 0.907 0.659 0.900 0.829 0.821 0.974 0.969 7.06 (4.96)
NSA 0.874 0.898 0.717 0.967 0.721 0.926 0.979 0.967 2.88 (3.36)
TabNSA 0.859 0.908 0.690 0.973 0.898 0.918 0.984 0.959 1.62 (0.744)
  • †Results for TabMixer correspond to the default model.
TabNSA’s high AUC on CB (0.898) stems from its sparse attention, which mitigates overfitting in small datasets. For DS (0.690), its lower performance may reflect its complex feature interactions, potentially requiring further tuning of NSA’s selection block.
5.5 Multi-class Classification
Table 4 presents a performance comparison of TabNSA and other neural network models on two multi-class classification datasets. Since multi-class classification tasks often suffer from class imbalance, we used macro-F1 as the primary evaluation metric during training to ensure a balanced assessment across all classes, following [30]. The results show that TabNSA consistently outperforms all models on both datasets.
Table 4: Comparison of different methods on SG and FO datasets.
@@ 5.5 Multi-class Classification
MLP TabTrans † TabNet TabNSA